North American prairie dogs live in underground burrows with several exits. They usually build a mount over one of the exits, which causes a draft to pass over the mound. How does this allow for ventilation of the burrow? [Hint: keep it simple – assume that all exits are at the same level and that there are no drafts over the other exits.]



1
Expert's answer
2021-03-14T19:12:23-0400

Explanation


  • Refer to the sketch attached


  • Bernoulli's principle is employed by these animals here in this act.
  • When we think the speed of air just above the other exists (one of them shown as A here) be zero & some speed just above the mounting area the hypothetical streamlines of air can be drawn as shown.


  • Since the speed of air near other exists is zero the pressure of air can be thought to have a higher value than that just above the mount where the speed of air is considerably higher.


  • This can be thought according to the Bernoulli equation (with potential energy part neglected), "\small k= P_A+\frac{1}{2}\rho v_A^2=P_B+\frac{1}{2}\rho v_B^2"


  • This results in a generation of a pressure difference across the burrow from A to B which facilitates some airflow through the burrow providing good ventilation to the burrow.

"\qquad\qquad\qquad\small P_a-P_B=\Delta P=\frac{1}{2}\rho v_B^2"